hassan and asghar had 80 marbles.hassan had four times as many as asghar had.how many marbles did each of hassan and asghar have?
Answer by dkppathak(439)

LET HASSAN HAVE MARBELS x
ASGHAR HAVE MARBELS Y
AS PER CONDITION GIVEN
X=4Y i
AND
X+Y =80 ii
PUTTING VALUE OF EQN i IN EQN ii
4Y+Y=80
5Y=80
Y=80/5=16
HASSAN HAVE 16X4 =64
ASGHAR HAVE =16